Circular No. 7010

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 M. H. Finger, Universities Space Research Association; S. W. Die-

ters, University of Alabama, Huntsville; and R. B. Wilson, Marshall Space Flight Center, NASA, report for the BATSE/Compton Observatory team: “A quasiperiodic oscillation (QPO) in the flux of this transient x-ray source is seen in power spectra of the BATSE 20–50-keV discriminator rates. On Sept. 9.0 UT, the QPO was centered on 271 ± 2 mHz, and it had a FWHM of 30 ± 5 mHz and a fractional rms of 19.2 ± 2.5 percent. The QPO was first detected on Sept. 7.45 near 55 mHz and has been steadily increasing in frequency, reaching 350 mHz by Sept. 9.3. The QPO appears to have reached the Nyquist frequency (488 mHz) of our discriminator rates before the end of Sept. 9, and is no longer detectable in them.”

C. A. Wilson, B. A. Harmon, W. S. Paciesas, and M. L. McCollough, Marshall Space Flight Center, NASA, report for the BATSE team: “The xray transient XTE J1550–564 (IAUC 7008) is being detected with BATSE in the band 20–100 keV. The outburst was first visible between Sept. 6.9 and 7.0 UT. Average 20–100-keV flux levels are (in 10^{−8} erg cm^{−2} s^{−1}): Sept. 6.0–7.0, < 0.1; 7.0–8.0, 1.26 ± 0.08, 8.0–8.6, 1.90 ± 0.12; 8.6–9.0, 2.42 ± 0.11. The spectrum is adequately fit by a power law. On Sept. 7, it extends beyond 200 keV and softens significantly during the intensity rise. Average photon number indices are: Sept. 7.0–8.0, −2.11 ± 0.09, 8.0–8.6, −2.62 ± 0.11; 8.6–9.0, −2.73 ± 0.09.”

D. Campbell-Wilson, V. McIntyre, R. Hunstead, and A. Green, University of Sydney, report observations at 843 MHz with the Molonglo Observatory Synthesis Telescope (MOST): “We detect a likely radio counterpart to the x-ray transient XTE J1550–564 in images obtained on Sept. 9 and 10. The source is located at α = 15^{h}50^{m}59^{s}.1, δ = −56^{o}28′39′′ (equinox 2000.0), with estimated errors of 5′′ in each coordinate, coincident with the possible optical counterpart reported on IAUC 7009. Archival MOST images obtained on 1992 Apr. 10 and 1998 June 14 show no radio source at this position to 3σ levels of 4.5 and 3.9 mJy, respectively. Preliminary flux densities are 10 mJy on Sept. 9.18 and 16 mJy on Sept. 10.16 UT (mid-exposure) with estimated errors of 2.5 mJy; image quality in these short observations is affected by sidelobes of the strong supernova remnant G326.3–1.8 immediately to the northeast.”

1998 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green